40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3573     Entry time: Wed Sep 15 01:27:52 2010
Author: rana, steve, valera 
Type: Update 
Category: PSL 
Subject: FSS cables connected 

- connected the TTFSS cables (FSS fast goes directly to NPRO PZT for now)

- measured the reference cavity 21.5 MHz EOM drive to be 17.8 dBm

-  turned on the HV for the FSS phase correcting EOM (aka PC) drive

- connected and turned on the reference cavity temperature stabilization

- connected the RefCav TRANS PD

- fine tuned the RefCav REFL PD angle

ELOG V3.1.3-